OpenStudy (anonymous):

Write the quadratic -1/2(x-2/7)^2-37/294 in general form?

OpenStudy (anonymous):

-1/2x^2+2/7x-1/6? Is this correct?

OpenStudy (anonymous):

or -1/2x^2+2/7x-2601/19306?

OpenStudy (imtant):

it is -1/2x^2-61/294..

OpenStudy (anonymous):

isn't the the general form of a quadratic ax^2+bx+c? This is how I started it: \[-\frac{ 1 }{ 2 }(x-\frac{ 2 }{ 7 })^{2}-\frac{ 37 }{ 294 }\] \[-\frac{ 1 }{ 2 }(x - \frac{ 2 }{ 7 })(x - \frac{ 2 }{ 7 })-\frac{ 37 }{ 294 }\] \[-\frac{ 1 }{ 2 }(x^{2}-\frac{ 2 }{ 7 }x-\frac{ 2 }{ 7 }x+\frac{ 4 }{ 49 })-\frac{ 37 }{ 294 }\]

OpenStudy (anonymous):

\[\prod_{31\cot 1231\tan^{-1} 12312\sin^{-1} \leftarrow213\lim_{\iota123\sqrt{123} \rightarrow ?}}^{?}\]

OpenStudy (imtant):

Ya..you can also solve like that..

OpenStudy (anonymous):

That is the answer :-)